OpenStudy (anonymous):

solve (x-1)(x-2)/x-3 >_ 0 please help me so this step by step

OpenStudy (anonymous):

you want to know where it is positive, which his a synonym for \(>0\) it changes sign at the zero of each factor, namely at \(1,2,3\) so you need to consider four intervals \[(-\infty,1), (1,2),(2,3),(3,\infty)\]

OpenStudy (anonymous):

test the expression at any point in any one of these intervals, it doesn't matter what point you pick or what interval you are in. you could pick say \(x=0\) or \(x=4\) or anything i will pick \(x=0\)

OpenStudy (anonymous):

if we replace \(x\) by \(0\) we get \[\frac{-1\times (-2)}{-3}=-\frac{2}{3}\] which is negative.

OpenStudy (anonymous):

than means it is negative on the interval \((-\infty,1)\) then positive on the interval \((1,2)\) then negative on the interval \((2,3)\) and finally positive on the interval \((3,\infty)\)

OpenStudy (anonymous):

you want to know where it is positive, but notice you know where it is positive, negative, zero, and undefined. your solution is \[(1,2)\cup (3,\infty)\]

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Thank you!